Two people push a shopping cart with equal force. The cart is accelerating 0.75 m/s2, and the force each person is exerting is 34 N. There is a 0.5 N force of friction, an 882 N force of gravity, and an 882 N normal force. What is the mass of the shopping cart?

1
Expert's answer
2021-09-30T18:27:56-0400

The mass can be found by Newton's second law:


"ma=F-\mu N,\\\space\\
m=\frac{F-\mu N}{a}=\frac{34-0.5}{0.75}=44.7\text{ kg}."